The sum of two numbers should be -5 but at the same time their product must be 6
i dont get it...
i think it means
when you multiply 2 numbers it = 6
when you add them they -5
What can those numbers be?

Answer by ewatrrr(24785)


Hi
sum of two numbers is -5 and their product is 6 (-3 and -2 would work)
finding algebraically
Let x and (-5-x) represent the numbers
x(-5-x) = 6
-5x - x^2 = 6
x^2 + 5x + 6 = 0
(x+2)(x+3) = 0
(x+2)=0 x = -2 and (-5-(-2) = -3
(x+3)=0 x = -3 and (-5-(-3) = -2
CHECKING our Answer***
-3(-2) = 6